OpenStudy (anonymous):

george is 7 and his mom is 37. in how many years will his mother be 3 times as old as he is.?

OpenStudy (anonymous):

8 years

OpenStudy (anonymous):

Let the mothers age be 3x, and the son be x. The difference in their ages will always be the same. So, 3x - x = 30 x = 15. Checking, 15-7 = 8 37+8 = 45

OpenStudy (anonymous):

The age of his mom now + x years will be equal to 3 times his age now + x years 3 * (7 + x) = (37 + x) 21 + 3x = 37 + x 2x = 16 x = 8

OpenStudy (anonymous):

The answer is 8, they are both correct. My explanation is correct too but I have not considered 'x' to be the current age.

OpenStudy (anonymous):

thank u